OpenStudy (anonymous):

If f(x) = 3x + 2 and g(x) = 2x – 2, what is (f – g)(x)? x + 4 x – 2 x 5x – 2 x – 4

OpenStudy (jacob902):

example If F(x)=3x-2 and g(x)= 2x-1, then g(f(x)) is equal to what?

OpenStudy (anonymous):

I need help I don't know how to do this

OpenStudy (jacob902):

g(f(x)) means that the function g is composed at the function f. You need to plug the value of f(x) into the function g and simplify. g(x) = 2x - 1 Replace x with the value of f(x) which is 3x - 2 and simplify. g(f(x)) = 2x - 1 g(3x - 2) = 2(3x - 2) - 1 g(3x - 2) = 6x - 4 - 1 g(3x - 2) = 6x - 5 So, g(f(x)) = 6x - 5

OpenStudy (jacob902):

theres an example dont want to give it out

OpenStudy (jacob902):

without learning

OpenStudy (jacob902):

nothing

OpenStudy (anonymous):

so what is it? can u take me through the steps

OpenStudy (jacob902):

If f(x)=3x^2+1 and g(x)=2x+2 find a) g^-1(x), the inverse of g(x) b) (fog^-1)(2)

OpenStudy (jacob902):

If g(x) = 2x + 2, this means it takes an input value, doubles it, and then adds 2. The inverse will do the opposite, in the opposite order. So the inverse will first do the opposite of adding 2, so it subtracts 2, or x - 2. Then it does the opposite of multiplying, by dividing by 2, so (x-2) / 2. So g-1(x)= (x-2) / 2. To find (fog-1)(2), you start by putting 2 into g inverse for x . So g-1(2) = (2-2) /2 , so zero. Then, we put 0 into f, so f(0) = 3(0)2 + 1, so (fog-1)(2) = 1. Hope that helped.

OpenStudy (anonymous):

great explanation :)

OpenStudy (anonymous):

but what is the answer

OpenStudy (anonymous):

\[(f-g)(x)=f(x)-g(x)\] \[\therefore (f-g)(x)=(3x-2)-(2x-2)\] \[=3x-2-2x+2\] \[=3x-2x+2-2\] \[=x+0\] \[\therefore (f-g)(x)=x\]

OpenStudy (anonymous):

so it is x=0 right

OpenStudy (anonymous):

answer is x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

cheers!

OpenStudy (anonymous):

If f(x) = 2x2 + 3x and g(x) = x – 2, what is (f + g)(2)? 16 14 12 10 8

OpenStudy (anonymous):

it is 14 right